# Beyond one-loop: higher-order effects on Gross-Neveu-Yukawa tensorial criticality

### v2

@@ I Introduction
One of us has with Mandal recently proposed a unified description of insulating and superconducting order parameters that open the relativistic mass-like gap in the fermion spectrum [25]. In graphene for example, this results in the above order parameters becoming the components of a symmetric traceless tensor under the group SO(8). More generally, for $N$ two-component Dirac fermions, the general order parameter is a second-rank irreducible symmetric tensor under SO($2N$). Mean-field analysis of the Gross-Neveu model, which embodies such a symmetry, suggests the existence of a phase transition at a novel critical point where SO($2N$) symmetry is not preserved as it would normally be, but becomes spontaneously broken to SO($N$)$\times$×SO($N$). This possibility has been demonstrated for the specific extension of the Gross-Neveu model to a large number of fermion flavors [26].
The GNY field theory for the symmetric traceless rank-two tensor field generically features two types of quartic self-interactions [27, 28]: ‘double-trace’ and ‘single-trace’ quartic terms. While the double-trace quartic interaction appears in the field theories for the conventional vector-like order parameters, the single-trace quartic interaction emerges specifically from the tensor structure of the order parameters. The interplay of these two self-interactions during the renormalization group flow determines the ultimate nature of the phase transition in the theory. One-loop calculation suggests that the critical point and the concomitant continuous phase transition do not exist in the Gross-Neveu model (equivalent to GNY theory with $N_{f}=1$) with $2N\geq 4$, and that the SO($2N$) to SO($N$)$\times$×SO($N$) transition is weakly first order [26, 28]. The situation changes however if one introduces an independent number $N_{f}>1$ of fermion flavors [28]: for a fixed $N$, increasing such an independent $N_{f}$ inevitably yields a stable fixed point of the RG flow. The one-loop critical value of $N_{f}$ for a given $N$ has been established in the previous work, suggesting that the Gross-Neveu value of $N_{f}=1$ lies firmly in the regime of the fluctuation-induced first-order transition for $N=2$ and $N=4$, for example [28].
In this study we investigate the higher-order corrections to our results by analyzing the two-loop RG flow equations. We find that the overall dependence of the nature of the phase transition on $N$ and $N_{f}$ identified in the previous one-loop computations [28] remains robust in our two-loop results, allowing us to extract the $O(\epsilon)$ corrections to the critical values of $N_{f}$. Importantly, we find that $N_{f,c1}$ is increased while $N_{f,c2}$ is decreased by the two-loop corrections, narrowing the intermediate region where the critical fixed point exists in the unstable region of the coupling space. The decreased value of $N_{f,c2}$ indicates that fewer fermions may be required for a continuous phase transitions at a given $N$ than previously taught. For the specific case of SO(8) symmetry relevant to the spin-1/2 fermions in graphene, for example, we observe approximately 34% reduction from the one-loop results. Furthermore, we calculate the critical exponents at the next-leading order, providing a more comprehensive understanding of these quantum phase transitions.
@@ II Model
[…]
where $\psi$ is $d_{\gamma}NN_{f}$ component Majorana fermion, $\bar{\psi}\equiv\psi^{\intercal}\gamma_{0}$, $S=\sum_{i=1}^{N_{s}}\varphi_{i}\mathbb{S}^{i}$, and $\varphi_{i}$ is order parameter. Here, $S$ transforms as $S\rightarrow OSO^{\intercal}$ where $O\in\text{SO}(N)$, and $\mathbb{S}^{i}$ is $N\times N$ symmetric traceless matrix with $\text{Tr}[\mathbb{S}^{i}\mathbb{S}^{j}]=\bar{N}_{\text{Tr}}\delta_{ij}$ ($i,j=1,\cdots,N_{s}=(N-1)(N+2)/2$). $\gamma_{a}$ ($a=0,\cdots D-1$) satisfies the Clifford algebra, $\{\gamma_{a},\gamma_{b}\}=2\delta_{ab}\mathbb{I}_{d_{\gamma}}$ where $d_{\gamma}=2^{\lfloor D/2\rfloor}$, and we assume $\gamma_{0}$ is antisymmetic. Note that we will set $d_{\gamma}=2$ at the end of the computation because we consider $D=2+1$ dimensions, so $\psi$ is $2NN_{f}$ component Majorana fermion. The parameters $\bar{g}$ and $\bar{\lambda}_{1,2}$ are bare coupling constants for the Yukawa interaction and quartic interactions, respectively, which have the engineering scaling dimension $4-D$.
The model features two types of self-interactions allowed by the symmetry: single-trace and double-trace quartic interactions. When $N\geq 4$, two quartic terms are mutually independent. The presence of two quartic interactions is responsible for a non-trivial RG flows and the possibility of the fluctuation-induced first-order transition in the theory [27, 30, 28]. We omit the $\text{Tr}[S^{3}]$ term, as such interactions cannot be generated by integrating out fermions. Consequently, the resulting Lagrangian exhibits an enhanced $\text{SO}(N)\times\mathbb{Z}_{2}\simeq O(N)$ symmetry, where the $\mathbb{Z}_{2}$ transformation corresponds to $S\rightarrow-S$ [28].
@@
where $\bar{\varphi}_{0}$ is the real-valued order parameter amplitude. In this case, SO($N$) symmetry is broken to SO($N-1$). The stability of the theory also demands that the following condition is satisfied:
@@ III Two-loop RG flow equations
We compute the two-loop order RG flow equations around $D=4$ space-time dimensions by using the dimensional regularization with the modified minimal subtraction ($\overline{\text{MS}}$) scheme [31, 29]. As a check of our computation we have also used RGBeta [32] which automates this calculation, and found the consistent results. The two-loop order RG flow-equations are given by
@@
and $\ell=\ln(\Lambda/\mu)$. Here, we set $\bar{r}=0$ and $d_{\gamma}=2$ to get the result for $D=3$ case. Note that replacing $N_{f}\rightarrow 2N_{f}$ one gets the flow-equations in $D=4$ because $d_{\gamma}=4$ in $D=4$. Also, one can obtain the flow-equations for $N=2,3$ by simply considering the flow of $\lambda=\lambda_{1}+(N/2)\lambda_{2}$, because, as we already mentioned, the two quartic ivariants are in this case proportional [28, 29].
@@ III.1 Critical $N$ without fermions
For example, this yields $N_{c1}\approx 2.077$ and $N_{c2}=3.535$ for $\epsilon=1$. This is consistent with a previous study on the antiferromagnetic RP model [27]. Therefore, in the absence of fermions, $N_{c1}$ and $N_{c2}$ both decrease in the two-loop order computation, with $N=2$ and $N=3$ theories still possesing critical fixed points [29], whereas $N\geq 4$ is not.
@@ III.2 Critical $N_{f}$ with fermions
Evidently, the two-loop order contribution reduces the value of $N_{f,c2}$ for given $N$ and $\epsilon$. Note that for the odd $N$ case, there is an additional prefactor $(N^{2}+3)/(N^{2}-1)$ (see Eq. 7), but this approaches 1 in the large $N$ limit. Therefore, we can use the same $N_{f,c2}$ for both even and odd values of $N$ in this limit.
@@ III.3 Universal quantities to two-loops
By substituting the fixed-point value into the above expression, the corresponding actual critical exponents can be obtained. For given $N$, the critical exponents in terms of $x=N_{f}-N_{f,c2}$ with $\epsilon=1$ are presented in Fig. 2. Furthermore, we can consider the mass gap ratio,
@@
Note that here $\mathcal{R}_{G}$ is given up to $\epsilon$-linear corrections because the one-loop computations give us $\epsilon^{0}$ order values [36, 37, 28]. We also present, in Fig. 3, a comparison among the one-loop and two-loop results and their fitted values shown in Eqs. (23)–(26) for each quantity at $N=8$ with $\epsilon=1$. Note that, as exemplified in the case of $N=8$, the functional forms used in Eqs. (23)–(26) provide good fits for various values of $N$ and $x$ as well.
@@
III.4 Honeycomb lattice with SO(8) symmetric Majorana fermions
@@ IV Three-loop corrections
Again, these are consistent with the previous study [27]. The series as usual does not appear to be convergent, and one needs a resummation [27]. From the Borel-Padé approximation, for example, for $\epsilon=1$, we can obtain $N_{c1}=2.387$ and $N_{c2}=3.595$.
@@ V Summary and Discussion
To further validate our findings, we have extended our analysis to include three-loop results for $N_{f,c2}$. Notably, the coefficient of $\epsilon^{2}$ exceeds that of $\epsilon$, highlighting a common artifact of the $\epsilon$-expansion. By employing Borel-Padé approximation, we confirm that the three-loop order $N_{f,c2}$ remains lower than the one-loop result, reinforcing the trend observed in our two-loop calculations. However, comprehensive validation would ultimately require comparison with numerical simulations and other methods.
Although we have only a few low-order terms with coefficients that suggest a non-convergent series, necessitating resummation methods, the alternating signs of these coefficients suggest that the epsilon expansion may eventually converge. Since the two-loop correction carries a negative sign, we can reasonably expect that the ultimate value converges to something smaller than the leading-order result.

### v3

# Beyond one-loop calculation: Higher-order effects on Gross-Neveu-Yukawa tensorial criticality

@@ I Introduction
[…]
In this study we investigate the higher-order corrections to our results by analyzing the two-loop RG flow equations. We find that the overall dependence of the nature of the phase transition on $N$ and $N_{f}$ identified in the previous one-loop computations [28] remains robust in our two-loop results, allowing us to extract the $O(\epsilon)$ corrections to the critical values of $N_{f}$. Importantly, we find that $N_{f,c1}$ is increased while $N_{f,c2}$ is decreased by the two-loop corrections, narrowing the intermediate region where the critical fixed point exists in the unstable region of the coupling space. The decreased value of $N_{f,c2}$ indicates that fewer fermions may be required for a continuous phase transitions at a given $N$ than previously thought. For the specific case of SO(8) symmetry relevant to the spin-1/2 fermions in graphene, for example, we observe approximately 34% reduction from the one-loop results. Furthermore, we calculate the critical exponents at the next-leading order, providing a more comprehensive understanding of these quantum phase transitions.
@@ II Model
$\quad\quad+\int d\tau d^{d}x\frac{1}{\bar{N}_{\text{Tr}}}\left[\frac{1}{2}\text{Tr}[(\partial_{\tau}S)^{2}+(\nabla S)^{2}+rS^{2}]+\frac{\bar{\lambda}_{1}}{4\bar{N}_{\text{Tr}}}(\text{Tr}[S^{2}])^{2}+\frac{\bar{\lambda}_{2}}{4}\text{Tr}[S^{4}]\right],$ (1)
where $\psi$ is $d_{\gamma}NN_{f}$ component Majorana fermion, $\bar{\psi}\equiv\psi^{\intercal}\gamma_{0}$, $S=\sum_{i=1}^{N_{s}}\varphi_{i}\mathbb{S}^{i}$, and $\varphi_{i}$ is order parameter. Here, $S$ transforms as $S\rightarrow OSO^{\intercal}$ where $O\in\text{SO}(N)$, and $\mathbb{S}^{i}$ is $N\times N$ symmetric traceless matrix with $\text{Tr}[\mathbb{S}^{i}\mathbb{S}^{j}]=\bar{N}_{\text{Tr}}\delta_{ij}$ ($i,j=1,\cdots,N_{s}=(N-1)(N+2)/2$). $\gamma_{a}$ ($a=0,\cdots D-1$) satisfies the Clifford algebra, $\{\gamma_{a},\gamma_{b}\}=2\delta_{ab}\mathbb{I}_{d_{\gamma}}$ where $d_{\gamma}=2^{\lfloor D/2\rfloor}$, and we assume $\gamma_{0}$ is antisymmetic. Note that we will set $d_{\gamma}=2$ at the end of the computation because we consider $D=2+1$ dimensions, so $\psi$ is $2NN_{f}$ component Majorana fermion. The parameters $\bar{g}$ and $\bar{\lambda}_{1,2}$ are bare coupling constants for the Yukawa interaction and quartic interactions, respectively, which have the engineering scaling dimension 4D4-D.
The model features two types of self-interactions allowed by the symmetry: single-trace and double-trace quartic interactions. When $N\geq 4$, two quartic terms are mutually independent. The presence of two quartic interactions is responsible for a non-trivial RG flows and the possibility of the fluctuation-induced first-order transition in the theory [27, 30, 28]. Note that the cubic term, $\text{Tr}[S^{3}]$, is absent, due to the additional inversion ($\mathbb{Z}_{2}$) symmetry that exists in the presence of fermions: $S\rightarrow-S$, $\psi\rightarrow i\gamma_{0}\gamma_{1}\psi$, $\partial_{2}\rightarrow-\partial_{2}$, for example, leaves the action invariant [28].
@@
where $\bar{\varphi}_{0}$ is the real-valued order parameter amplitude. In this case, SO($N$) symmetry is broken to SO(N1N-1). The stability of the theory also demands that the following condition is satisfied:
@@ III Two-loop RG flow equations
We compute the two-loop order RG flow equations around $D=4$ space-time dimensions by using the dimensional regularization with the modified minimal subtraction ($\overline{\text{MS}}$) scheme [31, 29]. As a check of our computation we have also used RGBeta [32, 33] which automates this calculation, and found the consistent results. The two-loop order RG flow-equations are given by
@@
and $\ell=\ln(\Lambda/\mu)$. Here, we set $\bar{r}=0$ and $d_{\gamma}=2$ to get the result for $D=3$ case. Note that by replacing $N_{f}\rightarrow 2N_{f}$ one gets the flow-equations in $D=4$ because $d_{\gamma}=2$ in $D=3$ but $d_{\gamma}=4$ in $D=4$. Also, one can obtain the flow-equations for $N=2,3$ by simply considering the flow of $\lambda=\lambda_{1}+(N/2)\lambda_{2}$, because, as we already mentioned, the two quartic ivariants are in this case proportional [28, 29].
@@ III.1 Critical $N$ without fermions
For example, this yields $N_{c1}\approx 2.077$ and $N_{c2}=3.535$ for $\epsilon=1$. This is consistent with a previous study on the antiferromagnetic RP model [27]. Therefore, in the absence of fermions, $N_{c1}$ and $N_{c2}$ both decrease in the two-loop order computation, with $N=2$ and $N=3$ theories still possessing critical fixed points [29], whereas $N\geq 4$ is not.
@@ III.2 Critical $N_{f}$ with fermions
We present $N_{f,c1}/N$ and $N_{f,c2}/N$ in terms of $\epsilon$ in $N\rightarrow\infty$ limit in Fig. 1. Evidently, the two-loop order contribution reduces the value of $N_{f,c2}$ for given $N$ and $\epsilon$. Note that for the odd $N$ case, there is an additional prefactor $(N^{2}+3)/(N^{2}-1)$ (see Eq. 7), but this approaches 1 in the large $N$ limit. Therefore, we can use the same $N_{f,c2}$ for both even and odd values of $N$ in this limit.
@@ III.3 Universal quantities to two-loops
By substituting the fixed-point values into the above expressions the corresponding actual critical exponents can be obtained. For given $N$, the critical exponents in terms of $x=N_{f}-N_{f,c2}$ with $\epsilon=1$ are presented in Fig. 2. Furthermore, we can consider the mass gap ratio,
@@
Note that here $\mathcal{R}_{G}$ is given up to $\epsilon$-linear corrections because the one-loop computation gives us $\epsilon^{0}$-order values [8, 37, 28]. We also present, in Fig. 3, a comparison among the one-loop and two-loop results and their fitted values shown in Eqs. (23)–(26) for each quantity at $N=8$ with $\epsilon=1$. Note that, as exemplified in the case of $N=8$, the functional forms used in Eqs. (23)–(26) provide good fits for various values of $N$ and $x$ as well.
@@
III.4 Honeycomb lattice with SO(8)-symmetric Majorana fermions
@@ IV Three-loop corrections
Again, these are consistent with the previous study [27]. The series as usual does not appear to be convergent, and one needs a resummation [27]. From the Borel-Padé approximation, which is a combination of the Borel summation and Padé approximation [38, 39, 2] (for details, see Appendix C), for example, for $\epsilon=1$, we can obtain $N_{c1}=2.387$ and $N_{c2}=3.595$.
@@ V Summary and Discussion
To further validate our findings, we have extended our analysis to include three-loop results for $N_{f,c2}$. Notably, the coefficient of $\epsilon^{2}$ exceeds that of $\epsilon$, highlighting a common feature of the $\epsilon$-expansion. By employing Borel-Padé approximation, we confirm that the three-loop order $N_{f,c2}$ remains lower than the one-loop result, reinforcing the trend observed in our two-loop calculations. However, comprehensive validation would ultimately require comparison with numerical simulations and other methods.
Although we have only a few low-order terms with coefficients that suggest a non-convergent series, necessitating resummation methods, the alternating signs of these coefficients suggest that the epsilon expansion may be asymptotic. Since the two-loop correction to $N_{f,c2}$ is negative we expect that its ultimate value is still smaller than the leading-order result.
@@ Appendix B Three loop contributions in critical exponenets
Appendix C Borel-Padé approximation
In this section, we will introduce the Borel-Padé approximation. The Borel-Padé approximation is a combination of the Borel summation and Padé approximation [38, 39, 2]. The procedure is as follow: As an example, let us consider the following expansion in terms of $\epsilon$ up to $\epsilon^{2}$ order.
$f(\epsilon)={}$ $a_{0}+a_{1}\epsilon+a_{2}\epsilon^{2}+\mathcal{O}(\epsilon^{3}).$ (51)
First, the Borel transformation of $f(\epsilon)$ is given by
$\mathcal{B}[f](\epsilon)={}$ $\sum_{n=0}^{\infty}\frac{a_{n}}{n!}\epsilon^{n}$
$={}$ $a_{0}+a_{1}\epsilon+\frac{a_{2}}{2}\epsilon^{2}+\mathcal{O}(\epsilon^{3}).$ (52)
Then, we apply the Padé approximation into Eq. (52). In particular, we will consider the Padé approximation of order $[1/1]$, here. It is given by
$\mathcal{P}_{[1/1]}[\mathcal{B}[f](\epsilon)]={}$ $\frac{\bar{a}_{0}+\bar{a}_{1}\epsilon}{1+\bar{b}_{1}\epsilon},$ (53)
where
$\bar{a}_{0}={}a_{0},\quad\bar{a}_{1}=a_{1}-\frac{a_{0}a_{2}}{2a_{1}},\quad\bar{b}_{1}={}-\frac{a_{2}}{2a_{1}}.$ (54)
Finally, we get the Borel-Padé approximation for a given expression by performing the Borel integral transformation,
$\tilde{f}(\epsilon)={}$ $\int_{0}^{\infty}e^{-t}\mathcal{P}_{[1/1]}[\mathcal{B}[f](t\epsilon)]dt.$ (55)